please help me..- find the slope and the y-intercept of the graph of each equation. okay so the problem is x - y = -2. plz help me i dont understand very well:(
Answer by scott8148(6628)

the slope-intercept form of the equation is ___ y = mx + b, where m is the slope and b is the y-intercept

solving an equation for y puts it in slope-intercept form

x - y = -2 ___ -y = -2 - x ___ y = x + 2

so the slope is 1, and the y-intercept is 2